OpenStudy (anonymous):

Need help with solving with substitution and elimination. Tell whether the system has one solution, infinitely many solutions, or no solution. Solve using substitution: 13. –x + y = -13 3x – y = 19 15. 1/3y = 7/3x + 5/3 x – 3y = 5 17. 3x + y = -13 -2x +5y = -54 Solve using elimination: Tell whether the system has one solution, infinitely many solutions, or no solution. 19. x + 2y = 23 5x + 10y = 55 21. 5x + 4y = -83 3x - 3y = -12 23. 4x + y = 21 -2x + 6y = 9 Even if you can help by explaining it in simplest terms, that would be helpful.

OpenStudy (solomonzelman):

13. add the second equation to the first. solve for y, and plug in then for x. SHOW YOUR STEPS HERE! ( we are helping not giving answers on this website )

OpenStudy (lovelyharmonics):

13: y= x-13 now plug ^ equation into 3x-y=19

OpenStudy (lovelyharmonics):

as in, since y is equal to x-13 youre going to take x-13 and plug it into the other equation to replace y... so 3x- (put it here) =19

OpenStudy (texaschic101):

SUBSTITUTION : 13. -x + y = -13 y = x - 13 now sub x - 13 in for y in the other equation 3x - y = 19 3x - (x - 13) = 19 -- distribute through the parenthesis 3x - x + 13 = 19 -- subtract 13 from both sides 3x - x = 19 - 13 -- combine like terms 2x = 6 -- divide both sides by 2 x = 3 now sub 3 in for x in either equation to find y y = x - 13 y = 3 - 13 y = -10 check your answers... -x + y = -13 -3 + (-10) = -13 -3 - 10 = -13 -13 = -13 (correct) SOLUTION : x = 3 and y = -10 (ONE SOLUTION) ============================ ELIMINATION : 19. x + 2y = 23 5x + 10y = 55 (this can be reduced to x + 2y = 11) now we have : x + 2y = 23 x + 2y = 11 Same slope, different y intercepts, ...that means the lines are parallel and do not intersect, therefore, there is NO SOLUTION. However, if you want to see it worked out... x + 2y = 23 --->(-5) 5x + 10y = 55 --------------- -5x - 10y = - 115 (result of multiplying by -5) 5x + 10y = 55 ---------------add 0 + 0 = - 60 0 = 60 (incorrect) They will not intersect, NO SOLUTION ============================ There are two of the problems, one by substitution, and the other by elimination. Any questions, just ask :)

OpenStudy (anonymous):

Thank you so much for the help. I understand it better now!